OpenStudy (anonymous):

A line goes through the points (1, 2) and (5, 10). Which of the following is the equation of the line?

OpenStudy (anonymous):

Option please @Jose_garcia1007

OpenStudy (anonymous):

\[eq. ~of~line~thro'~two~points~is~y-y1=\frac{ y2-y1 }{x2-x1 }\left( x-x1 \right)\]

OpenStudy (texaschic101):

oh....thats good....I never thought of putting the slope formula into the point slope formula...that makes sense